# Cognitio Emergens: Agency, Dimensions, and Dynamics in Human–AI Knowledge Co-Creation

### v1

@@ 3.2.4 Dimensional Interdependencies and Capability Signatures
These six dimensions, while conceptually distinct, interact in complex, often synergistic ways within actual human-AI partnerships. Advancements in one dimension frequently enable or constrain developments in others. For example, strong Interpretive Intelligence typically facilitates deeper Divergent Intelligence by making novel AI proposals more accessible and trustworthy for human researchers. Conversely, high Connective Intelligence might remain underutilized without sufficient Synthesis Intelligence to integrate the identified connections into meaningful frameworks.
The uneven development of these dimensions across partnerships creates distinctive “capability signatures” that characterize their specific strengths and weaknesses at a given time. Some teams might demonstrate advanced Connective and Synthesis Intelligence (excelling at integrating existing knowledge) while showing limited development in Divergent or Anticipatory Intelligence (struggling to generate novel ideas or explore futures). Others might excel in Divergent Intelligence while struggling with Interpretive Intelligence, leading to innovative but poorly understood outputs.
These capability signatures are not merely descriptive but also diagnostic, revealing both current strengths and potential developmental pathways or imbalances. Visualizing these signatures (as discussed further in Section 4) provides a powerful tool for understanding partnership evolution, identifying strategic intervention points, and tailoring development efforts to specific research goals and contexts.
@@ 3.3 Partnership Dynamics
While Epistemic Dimensions describe specific capabilities that emerge within partnerships, Partnership Dynamics—the third framework component—address the evolutionary forces that shape how these capabilities develop over time. These dynamics explain why certain partnerships flourish while others stagnate, and why some achieve breakthrough innovations while others remain limited to incremental improvements. Unlike the more structural Agency Configurations or capability-focused Epistemic Dimensions, these dynamics capture the temporal, evolutionary aspects of human-AI collaboration. By understanding these forces, research teams can more effectively navigate developmental challenges, harness generative potential, and mitigate emergent risks as their partnerships mature.
Partnership Dynamics are organized into three categories: Generative Dynamics (creating transformative potential), Balancing Dynamics (managing tensions), and Risk Dynamics (introducing vulnerabilities). Together, these dynamics constitute a complex adaptive system that shapes how partnerships develop and what they ultimately achieve.
@@ 3.3.1 Generative Dynamics
Generative Dynamics are catalytic forces that drive transformative innovation within human-AI partnerships, enabling radical research breakthroughs and paradigm shifts beyond what either humans or AI could achieve independently.
Transformative Potential. Transformative Potential refers to the partnership’s capacity to create paradigm shifts, reshape knowledge structures, and enable breakthroughs that fundamentally alter research fields. This dynamic emerges when the partnership achieves sufficient computational scale, conceptual integration, and creative synergy to move beyond incremental improvements. It signifies the potential for the partnership to not just answer existing questions but to change the questions being asked.
Transformative Potential manifests when partnerships fundamentally rethink foundational assumptions in their field. It appears as the emergence of entirely new methodologies that neither human nor AI would have developed independently, the transcendence of traditional disciplinary boundaries, or the creation of entirely new knowledge domains. For instance, a partnership might develop a novel simulation technique that allows exploring previously inaccessible physical regimes, thereby revolutionizing a subfield.
This dynamic represents the partnership’s inherent capacity to catalyze epistemic revolutions, distinct from merely generating novel ideas (Divergent Intelligence) or connecting domains (Connective Intelligence). Understanding this potential helps teams recognize when they are positioned for breakthrough work.
Recursive Evolution. Recursive Evolution refers to the partnership’s capacity to continuously reshape its own foundations and processes through complex feedback loops. This dynamic emerges when the partnership develops sophisticated self-monitoring and adaptive capabilities, allowing progressive refinement and transformation of its own operations based on experience. The partnership learns not only about its research subject but also about how to conduct research more effectively.
When a human-AI partnership achieves advanced Recursive Evolution, it doesn’t simply operate within its parameters but actively transforms those parameters. The partnership might create documentation of its own evolution, develop new interaction protocols based on past successes or failures, or continuously refine its working relationship and shared understanding. The system doesn’t just solve problems but evolves how it approaches problem-solving itself.
The recursive nature of this dynamic means that advancements at higher levels can reshape capabilities at lower levels. This creates the potential for “evolutionary cascades,” where an advancement in one dimension at a higher agency level triggers transformations across multiple dimensions. Cultivating Recursive Evolution is key to long-term partnership sustainability and growth.
@@ 3.3.2 Balancing Dynamics
Balancing Dynamics represent forces that manage inherent tensions and trade-offs within human-AI partnerships. They enable sustainable advancement while maintaining epistemic integrity and meaningful human involvement.
Temporal Integration. Temporal Integration refers to the partnership’s capacity to operate across multiple timescales, recognize its own epistemic evolution over time, and modulate its operational tempo accordingly. This dynamic emerges when the partnership successfully integrates human historical intuition and contextual understanding with AI’s ability to process vast datasets and simulate multiple temporalities simultaneously. It involves managing the pace of discovery and integration effectively.
Advanced partnerships manifest Temporal Integration by developing awareness of their own knowledge evolution, perhaps by creating meta-level documentation of shifts in understanding over a project’s lifecycle. They consciously modulate their speed—sometimes deliberately slowing down for deep interpretation and validation, other times accelerating to rapidly explore breakthrough possibilities. The partnership can navigate between different temporal scales, from immediate data analysis to long-term projection and backcasting.
This dynamic represents the system’s awareness and management of its own temporal nature, distinct from merely detecting external shifts (Anticipatory Intelligence) or projecting futures. It enables the partnership to adapt its operational tempo according to epistemic needs rather than being driven solely by technological capabilities or external pressures, fostering more reflective and sustainable progress.
Epistemic Ambidexterity. Epistemic Ambidexterity refers to the partnership’s capacity to simultaneously manage potentially conflicting demands: balancing exploration (generating novel, potentially disruptive ideas) with exploitation (refining and validating existing knowledge), and balancing innovation potential with the need for human interpretive control. This dynamic emerges when the partnership develops mechanisms to modulate between exploratory and explanatory modes while maintaining epistemic integrity and ensuring meaningful human participation.
Here we positulate that organizations successful in human-AI knowledge creation often cultivate this dynamic. They balance pushing boundaries with creating conditions for feedback loops that gradually evolve organizational knowledge practices. Partnerships with strong Epistemic Ambidexterity develop structures and processes that preserve human interpretability without unduly sacrificing innovation potential. They establish feedback cycles that continuously adjust this balance based on context and goals.
This dynamic represents the delicate equilibrium between pushing epistemic boundaries and maintaining human understanding that characterizes the most productive and sustainable human-AI partnerships. It involves not just possessing divergent and interpretive capabilities, but skillfully calibrating between them in response to evolving research needs and contexts. Achieving this balance is crucial for harnessing AI’s power without losing human agency.
@@ 3.3.3 Risk Dynamics
Risk Dynamics represent forces that introduce vulnerabilities into human-AI partnerships. If not properly managed, these dynamics can undermine epistemic integrity, trust, and the overall value of the collaboration.
@@
Epistemic Alienation. Epistemic Alienation refers to a phenomenon where researchers feel dispossessed of interpretive ownership over knowledge they formally endorse. They may accept AI-generated outputs as valid but struggle to fully explain, contextualize, or defend them within their existing epistemic frameworks. This dynamic emerges when the gap between AI-generated outputs and human interpretive capacities exceeds a critical threshold, often linked to insufficient Interpretive Intelligence (Section 3.2.1).
@@
Unlike classic automation complacency (passive reliance) or system opacity problems (inability to trace causal chains), Epistemic Alienation denotes a deeper normative and existential loss. Researchers feel disconnected from the meaning and foundations of the knowledge, even if they understand its provenance. It represents a breakdown in the researcher’s relationship with the knowledge itself.
@@
Epistemic Alienation risk generally increases with agency configuration. In Directed Agency, alienation is minimal as humans maintain tight interpretive control. In Contributory Agency, partial alienation may emerge around specific AI-initiated contributions. In Partnership Agency, where knowledge production becomes truly emergent and potentially less traceable to individual contributions, the risk of profound alienation increases significantly.
@@
When researchers experience Epistemic Alienation, they may retreat from more advanced agency configurations to regain interpretive control, potentially sacrificing innovative potential for epistemic security. Managing this dynamic requires deliberate attention to interpretive practices, fostering high Interpretive Intelligence, and developing new explanatory frameworks that bridge human and AI epistemologies.
@@
Epistemic Closure. Epistemic Closure refers to the narrowing of perspective that can occur through reinforcing feedback loops between humans and AI systems. This dynamic emerges when partnerships begin to inadvertently reinforce existing biases (either human or algorithmic) and reduce interpretive diversity. It is particularly concerning when AI is trained on biased or partial datasets, or when human researchers interact with AI in ways that consistently confirm their prior commitments.
@@
Preliminary interviews with AI-augmented research teams reveal how partnerships risk epistemic closure when feedback loops filter out dissenting perspectives or novel interpretations. The partnership may inadvertently create an echo chamber that amplifies certain viewpoints while dismissing alternatives, gradually reducing the diversity of theoretical approaches considered or methodologies employed.
@@
This dynamic represents a significant risk to the transformative potential of human-AI partnerships, as true innovation often requires challenging established assumptions and considering diverse viewpoints. Managing Epistemic Closure requires deliberately introducing diverse perspectives (e.g., through team composition, varied data sources, or incorporating AI designed to play devil’s advocate) and establishing mechanisms for periodically questioning the partnership’s core assumptions and reinforcing loops.
@@ 3.3.4 Dynamics Interactions and Threshold Effects
Partnership Dynamics create complex interaction patterns that significantly influence how human-AI collaborations evolve over time, generating both virtuous and vicious cycles that shape development trajectories. For example, strong Epistemic Ambidexterity can mitigate Epistemic Alienation risk by ensuring human understanding keeps pace with innovation. Conversely, unchecked Transformative Potential without sufficient Epistemic Ambidexterity or Interpretive Intelligence may accelerate Epistemic Alienation.
Particularly significant are potential threshold effects that may occur when dynamics reach critical levels or interact synergistically. When multiple generative dynamics exceed certain thresholds simultaneously, qualitative transformations in knowledge production—true paradigm shifts—may become possible. For instance, when high levels of Recursive Evolution and Transformative Potential coincide within a partnership exhibiting strong Epistemic Ambidexterity, rapid transitions to stable Partnership Agency configurations might occur, enabling breakthrough innovations that reshape entire fields.
Understanding these dynamic interactions and potential thresholds helps research teams anticipate and navigate epistemic transformations more effectively, rather than experiencing them merely as unexpected disruptions. The framework provides conceptual tools for identifying potential vulnerabilities (e.g., risk of Epistemic Closure) and leveraging opportunities (e.g., harnessing Transformative Potential) as partnerships evolve.
@@ 3.4 Theoretical Implications
The Cognitio Emergens framework transforms how we understand knowledge creation in the age of advanced AI, with particularly profound implications for epistemology and agency. These primary implications, along with important secondary effects on methodology, organizations, and ethics, extend beyond specific research contexts to address fundamental questions about the nature of human-AI collaboration.
@@ 3.4.1 Epistemological Implications
The framework suggests a fundamental shift in epistemology. It points beyond both individual human cognition and purely social knowledge creation to a third epistemological space where knowledge emerges dynamically from human-AI partnerships. This represents a novel form of distributed cognition that transcends current epistemological categories, demanding new theories of knowing.
While Thomas Kuhn described how scientific paradigms shift through crisis and revolution (Kuhn & Hacking, 2012), this framework suggests how paradigms themselves might be systematically explored, evaluated, and perhaps even generated through mature human-AI partnerships possessing advanced dimensional capabilities (especially Divergent, Synthesis, and Axiological Intelligence). CE potentially offers a mechanism for more deliberate paradigm navigation and evolution.
The framework also highlights profound questions about the nature of knowledge itself. If knowledge traditionally requires not just production but also interpretive ownership and justification by human knowers, how do we conceptualize insights emerging from systems where human interpretive capacity may be outpaced by generative capability (Shanahan, 2025)? This challenges us to rethink what it means to “know” in a human-AI context.
@@ 3.4.2 Agency Implications
The framework reconceptualizes agency in knowledge creation as distributed, dynamic, and emergent, rather than singular, fixed, and located solely in the human (Holter & El-Assady, 2024; Watkins et al., 2025). Agency arises from the structured interaction between humans and AI, creating forms of distributed cognition that possess capabilities beyond those of any individual component. This contrasts with more agent-centric views (Silver & Sutton, 2025).
This perspective challenges traditional assumptions about epistemic authority and responsibility. As agency configurations evolve toward Partnership, the locus of effective epistemic authority shifts from individual humans to the human-AI system itself. This raises critical questions about accountability, attribution of discovery, and the definition of scientific contribution in an era of advanced AI, touching on governance issues highlighted in recent studies (Liu & Shen, 2025; Lou et al., 2025).
@@ 3.4.3 Methodological Implications
The framework suggests that research methodology itself is not static but can co-evolve with the human-AI partnership. Through dimensional development and agency transitions, research methods can be transformed, reflecting multi-loop learning processes (Argyris, 1976; Lou et al., 2025; Tosey et al., 2012). The most advanced partnerships don’t just apply existing methodologies more efficiently—they can create fundamentally new methodological approaches that transcend current disciplinary boundaries.
This implies a shift in how we evaluate research methods. Instead of viewing methods as fixed tools, we can see them as evolving capabilities of the partnership. This allows for richer questions: “In which dimensions (e.g., Synthesis, Anticipatory) has this partnership’s methodological approach evolved furthest? Where are the developmental imbalances? What novel methodologies might emerge if we strategically cultivate specific dimensions?”

### v2

@@ 3.2.4 Dimensional Interdependencies and Capability Signatures
These six dimensions, while conceptually distinct, interact in synergistic ways within partnerships. Advancements in one dimension enable or constrain developments in others. Strong Interpretive Intelligence facilitates deeper Divergent Intelligence by making novel AI proposals more accessible and trustworthy. Conversely, high Connective Intelligence might remain underutilized without sufficient Synthesis Intelligence to integrate identified connections into meaningful frameworks.
Uneven development across dimensions creates distinctive “capability signatures” characterizing specific strengths and weaknesses. Some teams might demonstrate advanced Connective and Synthesis Intelligence (excelling at integrating existing knowledge) while showing limited Divergent or Anticipatory Intelligence (struggling to generate novel ideas or explore futures). Others might excel in Divergent Intelligence while struggling with Interpretive Intelligence, leading to innovative but poorly understood outputs.
Capability signatures are diagnostic, revealing current strengths and potential developmental pathways or imbalances. Visualizing these signatures (Section 4) provides a tool for understanding partnership evolution, identifying strategic intervention points, and tailoring development efforts to specific research goals and contexts.
@@ 3.3 Partnership Dynamics
Epistemic Dimensions describe capabilities that emerge within partnerships. How these capabilities develop over time, however, depends on evolutionary forces captured by Partnership Dynamics—the third framework component. These dynamics explain why partnerships flourish or stagnate, why some achieve breakthrough innovations while others remain limited to incremental improvements. These dynamics capture temporal, evolutionary aspects of human-AI collaboration.
Partnership Dynamics organize into three categories: Generative Dynamics (creating transformative potential), Balancing Dynamics (managing tensions), and Risk Dynamics (introducing vulnerabilities).
@@ 3.3.1 Generative Dynamics
Generative Dynamics are catalytic forces driving transformative innovation within partnerships, creating radical research breakthroughs beyond what either humans or AI could achieve independently.
Transformative Potential. Transformative Potential refers to capacity to create paradigm shifts, reshape knowledge structures, and drive breakthroughs altering research fields. This dynamic emerges when partnerships achieve sufficient computational scale, conceptual integration, and creative synergy to move beyond incremental improvements.
Transformative Potential manifests when partnerships rethink foundational assumptions. It appears as emergence of new methodologies that neither human nor AI would develop independently, transcendence of traditional disciplinary boundaries, or creation of new knowledge domains. A partnership might develop a novel simulation technique allowing exploration of previously inaccessible physical regimes, revolutionizing a subfield.
This dynamic catalyzes epistemic revolutions, distinct from generating novel ideas (Divergent Intelligence) or connecting domains (Connective Intelligence).
Recursive Evolution. Recursive Evolution refers to capacity to continuously reshape foundations and processes through feedback loops. This dynamic emerges when partnerships develop self-monitoring and adaptive capabilities, allowing progressive refinement and transformation of operations based on experience. Partnerships learn not only about research subjects but also how to conduct research more effectively.
When partnerships achieve advanced Recursive Evolution, they actively transform their parameters rather than simply operating within them. Partnerships might create documentation of their evolution, develop new interaction protocols based on past successes or failures, or continuously refine working relationships and shared understanding. The system evolves how it approaches problem-solving itself.
The recursive nature of this dynamic means advancements at higher levels can reshape capabilities at lower levels, creating potential for “evolutionary cascades” where advancement in one dimension at a higher agency level triggers transformations across multiple dimensions.
@@ 3.3.2 Balancing Dynamics
Balancing Dynamics represent forces managing tensions and trade-offs within partnerships. They enable sustainable advancement while maintaining epistemic integrity and meaningful human involvement.
Temporal Integration. Temporal Integration refers to capacity to operate across multiple timescales, recognize epistemic evolution over time, and modulate operational tempo accordingly. This dynamic emerges when partnerships integrate human historical intuition and contextual understanding with AI’s ability to process vast datasets and simulate multiple temporalities simultaneously.
Advanced partnerships manifest Temporal Integration by developing awareness of knowledge evolution, creating meta-level documentation of shifts in understanding over a project’s lifecycle. They consciously modulate speed—sometimes deliberately slowing for deep interpretation and validation, other times accelerating to rapidly explore breakthrough possibilities. Partnerships navigate between temporal scales, from immediate data analysis to long-term projection and backcasting.
This dynamic governs awareness and management of temporal nature, distinct from detecting external shifts (Anticipatory Intelligence) or projecting futures. It allows partnerships to adapt operational tempo according to epistemic needs rather than being driven solely by technological capabilities or external pressures.
Epistemic Ambidexterity. Epistemic Ambidexterity refers to capacity to simultaneously manage conflicting demands: balancing exploration (generating novel, potentially disruptive ideas) with exploitation (refining and validating existing knowledge), and balancing innovation potential with need for human interpretive control. This dynamic emerges when partnerships develop mechanisms to modulate between exploratory and explanatory modes while maintaining epistemic integrity and ensuring meaningful human participation.
Organizations successful in human-AI knowledge creation often cultivate this dynamic. They balance pushing boundaries with creating conditions for feedback loops that gradually evolve organizational knowledge practices. Partnerships with strong Epistemic Ambidexterity develop structures and processes preserving human interpretability without unduly sacrificing innovation potential. They establish feedback cycles that continuously adjust this balance based on context and goals.
This dynamic maintains equilibrium between pushing epistemic boundaries and preserving human understanding. It involves skillfully calibrating between divergent and interpretive capabilities in response to evolving research needs and contexts.
@@ 3.3.3 Risk Dynamics
Risk Dynamics represent forces introducing vulnerabilities into partnerships. If not managed, these dynamics can undermine epistemic integrity, trust, and collaboration value.
@@
Epistemic Alienation. Epistemic Alienation refers to where researchers feel dispossessed of interpretive ownership over knowledge they formally endorse. They may accept AI-generated outputs as valid but struggle to fully explain, contextualize, or defend them within existing epistemic frameworks. This dynamic emerges when the gap between AI-generated outputs and human interpretive capacities exceeds a critical threshold, often linked to insufficient Interpretive Intelligence (Section 3.2.1).
@@
Unlike automation complacency (passive reliance) or system opacity problems (inability to trace causal chains), Epistemic Alienation denotes deeper normative and existential loss. Researchers feel disconnected from the meaning and foundations of knowledge, even if they understand its provenance.
@@
Epistemic Alienation risk increases with agency configuration. In Directed Agency, alienation is minimal as humans maintain tight interpretive control. In Contributory Agency, partial alienation may emerge around specific AI-initiated contributions. In Partnership Agency, where knowledge production becomes emergent and less traceable to individual contributions, alienation risk increases significantly.
@@
When researchers experience Epistemic Alienation, they may retreat from more advanced agency configurations to regain interpretive control, potentially sacrificing innovative potential for epistemic security. Managing this dynamic requires deliberate attention to interpretive practices, fostering high Interpretive Intelligence, and developing new explanatory frameworks bridging human and AI epistemologies.
@@
Epistemic Closure. Epistemic Closure refers to narrowing of perspective through reinforcing feedback loops between humans and AI systems. This dynamic emerges when partnerships inadvertently reinforce existing biases (either human or algorithmic) and reduce interpretive diversity. It is concerning when AI is trained on biased or partial datasets, or when researchers interact with AI in ways that consistently confirm their prior commitments.
@@
Preliminary interviews with AI-augmented research teams reveal how partnerships risk epistemic closure when feedback loops filter out dissenting perspectives or novel interpretations. Partnerships may inadvertently create an echo chamber amplifying certain viewpoints while dismissing alternatives, gradually reducing diversity of theoretical approaches considered or methodologies employed.
@@
This dynamic poses significant risk to transformative potential of partnerships, as innovation often requires challenging established assumptions and considering diverse viewpoints. Managing Epistemic Closure requires deliberately introducing diverse perspectives (through team composition, varied data sources, or incorporating AI designed to play devil’s advocate) and establishing mechanisms for periodically questioning core assumptions and reinforcing loops.
@@ 3.3.4 Dynamics Interactions and Threshold Effects
Partnership Dynamics create complex interaction patterns influencing how collaborations evolve over time, generating both virtuous and vicious cycles shaping development trajectories. Strong Epistemic Ambidexterity can mitigate Epistemic Alienation risk by ensuring human understanding keeps pace with innovation. Conversely, unchecked Transformative Potential without sufficient Epistemic Ambidexterity or Interpretive Intelligence may accelerate Epistemic Alienation.
Particularly significant are potential threshold effects occurring when dynamics reach critical levels or interact synergistically. When multiple generative dynamics exceed certain thresholds simultaneously, qualitative transformations in knowledge production—paradigm shifts—may become possible. When high levels of Recursive Evolution and Transformative Potential coincide within a partnership exhibiting strong Epistemic Ambidexterity, rapid transitions to stable Partnership Agency configurations might occur, enabling breakthrough innovations reshaping entire fields.
Understanding these dynamic interactions and potential thresholds helps research teams anticipate and navigate epistemic transformations more effectively. The framework equips teams with conceptual tools for identifying potential vulnerabilities (risk of Epistemic Closure) and leveraging opportunities (harnessing Transformative Potential) as partnerships evolve.
@@ 3.4 Theoretical Implications
The Cognitio Emergens framework transforms how we understand knowledge creation in the age of advanced AI, with implications for epistemology and agency. These implications, along with effects on methodology, organizations, and ethics, extend beyond specific research contexts to address questions about the nature of human-AI collaboration.
@@ 3.4.1 Epistemological Implications
Most fundamentally, the framework suggests a shift in epistemology. It points beyond both individual human cognition and purely social knowledge creation to a third epistemological space where knowledge emerges dynamically from human-AI partnerships. This constitutes a form of distributed cognition that extends beyond current epistemological categories.
While Thomas Kuhn described how scientific paradigms shift through crisis and revolution (Kuhn and Hacking 2012), this framework suggests how paradigms themselves might be systematically explored, evaluated, and perhaps generated through mature partnerships possessing advanced dimensional capabilities (especially Divergent, Synthesis, and Axiological Intelligence).
The framework also highlights questions about the nature of knowledge itself. If knowledge traditionally requires production, interpretive ownership, and justification by human knowers, how do we conceptualize insights emerging from systems where human interpretive capacity may be outpaced by generative capability (Shanahan 2025)? This challenges us to rethink what it means to “know” in a human-AI context.
@@ 3.4.2 Agency Implications
The framework reconceptualizes agency in knowledge creation as distributed, dynamic, and emergent, rather than singular, fixed, and located solely in humans (Holter and El-Assady 2024; Watkins et al. 2025). Agency arises from structured interaction between humans and AI, creating forms of distributed cognition possessing capabilities beyond those of any individual component. This contrasts with agent-centric views (Silver and Sutton 2025).
This perspective challenges traditional assumptions about epistemic authority and responsibility. As agency configurations evolve toward Partnership, the locus of effective epistemic authority shifts from individual humans to the human-AI system itself. This raises questions about accountability, attribution of discovery, and definition of scientific contribution in an era of advanced AI, touching on governance issues highlighted in recent studies (Lou et al. 2025; Liu and Shen 2025).
@@ 3.4.3 Methodological Implications
The framework suggests that research methodology itself is not static but can co-evolve with partnerships. Through dimensional development and agency transitions, research methods can be transformed, reflecting multi-loop learning processes (Argyris 1976; Tosey, Visser, and Saunders 2012; Lou et al. 2025). Advanced partnerships don’t just apply existing methodologies more efficiently—they can create new methodological approaches that transcend current disciplinary boundaries.
This implies a shift in how we evaluate research methods. Instead of viewing methods as fixed tools, we can see them as evolving capabilities of partnerships. This allows richer questions: “In which dimensions (e.g., Synthesis, Anticipatory) has this partnership’s methodological approach evolved furthest? Where are developmental imbalances? What novel methodologies might emerge if we strategically cultivate specific dimensions?”